# Ashish Saxena 11    Tanvir Kaur 11    Kaushik Mondal 11

### v2

@@ Abstract
Besides being studied over static graphs heavily, the dispersion problem is also studied on dynamic graphs with $n$ nodes where at each discrete time step the graph is a connected sub-graph of the complete graph $K_{n}$. An optimal algorithm is provided assuming global communication and 1-hop visibility of the agents. How this problem pans out on Time-Varying Graphs (TVG) is kept as an open question in the literature. In this work, we study this problem on TVG considering $k\geq 1$ agents where at each discrete time step the adversary can remove at most one edge keeping the underlying graph connected. We have the following main results considering all agents start from a rooted initial configuration. Global communication and 1-hop visibility are must to solve dispersion for $pn$ ($p\geq 1$) co-located agents on a TVG even if agents have unlimited memory and knowledge of $n$. We provide an algorithm that disperses $n+1$ agents on TVG by dropping both the assumptions of global communication and 1-hop visibility using $O(\log n)$ memory per agent. We extend this algorithm to solve dispersion with $pn+1$ agents with the same model assumptions.
@@
Dispersion Mobile agents Distributed algorithms Time-Varying Graph Deterministic algorithms
@@ 1 Introduction
The dispersion problem, introduced in [3] on static graphs, involves the coordination of $k\leq n$ agents on the graph of size $n$ to reach a configuration with the constraint that at most one agent can be present at any node. This problem generally applies to real-world scenarios in which agents must coordinate and share resources at various places. The goal is to minimize the total cost of solving the problem while ensuring that the cost of agents moving around on the graph is much lower than the cost of having more than one agent share a resource. The dispersion problem is also connected to three other problems: (i) scattering or uniform deployment of agents on a graph [8], (ii) exploration of agents on a graph [9], and (iii) load balancing on a graph [6]. The scattering problem involves coordinating and moving $k\leq n$ agents on a given graph with $n$ nodes so that they are uniformly distributed over the graph. When $k=n$, the scattering problem is equivalent to the dispersion problem. The problem of collaborative exploration requires agents to visit all the nodes of a given graph while coordinating with one another. This problem is related to dispersion since if $n$ agents need to disperse, they must explore the graph. Load balancing on graphs is a problem where resources need to be evenly distributed as much as possible among nodes on a graph in the least possible number of rounds. Typically, nodes control the allocation of resources, and there are constraints on how much load can pass through an edge in a round. This problem is similar to dispersion if we consider agents as resources.
In [21], Molla et al. examine the dispersion problem for the first time for any value of $k\geq 1$. In this work, they have given the following definition of the dispersion.
Definition 1
Let $k$ agents be present at $n$ node graph $G$. Each node should have at most $\lceil k/n\rceil$ agents in the final configuration.
There are several works on dispersion [3, 13, 14, 22, 17, 23, 21], and most of them deal with static graphs. The addition of dynamism to a network makes the dispersion problem more challenging as well as practical. Agents need to complete their tasks in environments that are also changing. In this work, we aim to solve the dispersion problem for any $k$ on dynamic graphs. Below, we discuss the model assumption and the problem definition.
1.1 Model and the problem
Time-varying graph model: The system is represented as a time-varying graph (TVG), denoted by $\mathcal{G}=(V,E,\mathbb{T},\rho)$, where, $V$ represents the set of nodes, $E$ represents a set of edges, $\mathbb{T}$ represents the temporal domain, which is defined to be $\mathbb{Z}^{+}$ as in this model we consider discrete time steps, and $\rho:E\times\mathbb{T}\rightarrow\{0,1\}$ tells whether a particular edge is available at a given time. The graph $G=(V,\,E)$ is the underlying static graph of the dynamic graph $\mathcal{G}$, also termed as a footprint of $\mathcal{G}$. We denote $\delta_{v}$ as the degree of node $v$ in the footprint $G$, and $\Delta$ as the maximum degree in the footprint $G$. The graph $G$ is undirected and unweighted. The nodes of $G$ are anonymous (i.e., they have no IDs). The graph $G$ is port-labelled, i.e., each incident edge of a node $v$ is assigned a locally unique port number in $\{0,1,2,3,\cdots,\delta_{v}-1\}$. Specifically, each undirected edge $(u,\,v)$ has two labels, denoted as its port numbers at $u$ and $v$. Port numbering is local, i.e., there is no relationship between the port numbers at $u$ and $v$. We consider there is an adversary that can remove edges from the footprint $G$, keeping $\mathcal{G}$ connected at each round.
@@ 1.1 Model and the problem
Agent: We consider $k$ agents to be present initially at nodes of the graph $G$. Each agent has a unique identifier assigned from the range $[1,\,n^{c}]$, where $c$ is a constant. Each agent knows its ID. Agents are not aware of the values of $n$, $\Delta$, or $c$ unless these values are specifically mentioned. The agents are equipped with $O(\log n)$ memory. An agent residing at a node (say $v$) in round $t$ knows $\delta_{v}$ in the footprint $G$ and all associated ports corresponding to node $v$ in $G$; however, the agent does not understand if any incident edge of node $v$ is missing in $\mathcal{G}$ in round $t$. To be more precise, agent $a$ currently positioned at node $v$ at round $t$ does not know the value of $\rho(e_{v},t)$, where $e_{v}$ is an edge incident at node $v$. Such a model has been considered in [11, 12].
Communication model: There are two communication models in a distributed setting: local and global. In the local communication model [3, 13], an agent located at a particular graph node can communicate with other agents present at the same node. In contrast, the global communication model [10, 16] allows an agent at a graph node to communicate with any other agent in the graph, even if they are located at different nodes. In nature, this type of communication happens through the links of the graph. We use the local communication model in our algorithm.
Visibility model: In this work, we use two types of visibility models: zero-hop visibility and $1$-hop visibility. In the zero-hop visibility model [3, 14, 15], an agent located at a particular node (say $v\in\mathcal{G}$) can see agents present at node $v$ in round $r$, and port numbers associated with node $v$. In $1$-hop visibility [1, 20], an agent (say $a_{i}$) located at a particular node (say $v\in\mathcal{G}$) can see any node (along with agent(s) present at that node, if any) that is the neighbours of $v$ in $\mathcal{G}$. If the adversary deletes an edge associated with a node $u$ with port $p$ at the start of a round $r$, the agent(s) present there cannot understand that an edge has been deleted in the zero-hop visibility model. However, if the agent tries to move through that edge in this round, it cannot make the move. In the 1-hop visibility model, if such an edge is deleted in a round, then the agents present there can see all the neighbours of $v$ that are connected to $v$ via the edges with the port number $\neq p$ at $v$ because according to the 1-hop visibility model, the agents can see the sub-graph induced by the nodes that are 1-hop away in this round. Therefore, agents can also recognize that no node is connected through port $p$, and therefore, agents can understand that the edge corresponding to port $p$ at $v$ is deleted at the start of this round.
@@
The algorithm runs in synchronous rounds. In each round $t$, an agent $a_{i}$ performs one Communicate-Compute-Move (CCM) cycle as follows:
@@
  • Communicate: Agent $a_{i}$ at node $v_{i}$ can communicate with other agents $a_{j}$ present at the same node $v_{i}$ or present at any different node $v_{j}$, depending on the communication model used. The agent also understands whether it had a successful or an unsuccessful move in the last round.
@@
Now, we are ready to provide our problem definition.
Definition 2
$k$-Dispersion on TVG: Let $\mathcal{G}$ be a TVG. Let $k$ $(\geq 1)$ agents be at the nodes of the footprint $G$. At each round, the adversary can remove at most one edge from $G$, keeping $\mathcal{G}$ connected. The agents need to reposition themselves such that each node contains at most $\lceil k/n\rceil$ agents.
In our problem definition, we consider the adversary can delete at most one edge from footprint $G$ at each round and this restriction is already considered in several existing works [4, 5, 12].
ID Range IC Dynamism $\bm{k}$ Necessary Assumptions Sufficient Assumptions Algorithm
[16] [1, k] Scattered Arbitrary edge addition or deletion maintaining the graph connected at each round $k\leq n$ 1-hop visibility, global comm. 1-hop visibility, global comm. $\Theta(k)$ rounds, $\Theta(\log k)$ memory per agent
This work $[1,\,n^{c}]$, $c\in\mathbb{N}$ Co-located 1-TVG $k=n+1$ 0-hop visibility, local comm. $O(n^{4}\cdot\log n)$ rounds, $\Theta(\log n)$ memory per agent
$k=pn+1$, $p>1$ 0-hop visibility, local comm., knowledge of $n$ $O(n^{4}\cdot\max(\log n,\,\log p))$ rounds, $\Theta(\log n)$ memory per agent
Co-located 1-TVG $k=pn$, $p\geq 1$ 1-hop visibility, global comm.
Scattered $f$-TVG $k\geq 1$ 1-hop visibility, global comm. $O(k)$ rounds, $\Theta(\log n)$ memory per agent (An extension of [16])
Table 1: This table shows the results of dispersion on different dynamic graph models. In this table, IC represents the initial configuration, $1$-TVG means the adversary can delete at most one edge per round from footprint $G$ while maintaining $\mathcal{G}$ connected, and $f$-TVG means the adversary can delete at most $f$ edges per round from footprint $G$ while maintaining $\mathcal{G}$ connected.
1.2 Related work
The problem of dispersion for $k$ $(\leq n)$ agents has been extensively studied on static graphs [3, 13, 14, 22, 17, 23]. Here we focus on the literature of dispersion on dynamic graphs only. Recently, researchers in the field of distributed computing started studying dynamic graphs, where the topological changes are not sporadic or anomalous but rather inherent to the nature of the network. Such a dynamic graph model was developed by Kuhn et al. [18] in 2010. Dispersion on the dynamic graphs for $k\leq n$ agents is studied by Agarwalla et al. [1] and Kshemkalyani et al. [16]. In [1], the authors consider the following dynamic graph model on ring $\mathcal{R}$. In [1], the adversary can delete at most one edge from $\mathcal{R}$ at each round. Based on this dynamic graph model, they develop several deterministic algorithms for agents to achieve dispersion on a ring. In [16], the authors use a weaker dynamic graph model than [1]. In [1], the adversary can delete at most one edge per round from the ring, but in [16], the adversary can add or remove edges, keeping the dynamic graph connected at each round. The authors show that it is impossible to solve dispersion on the dynamic graph in the local communication model, even if $1$-hop visibility is available to the agents and each agent has unlimited memory. Also, it is impossible to solve dispersion on the dynamic graph in the global communication model without $1$-hop visibility, even with unlimited memory at each agent. They provided an asymptotically optimal $\Theta(k)$ rounds algorithm in the global communication model with $1$-hop visibility.
In [2], Casteigts et al. defined the new dynamic graph paradigm, which is known as TVG. In this model, the exploration [11], gathering [19], and black-hole search problem [7, 4, 5, 12] have been studied. The basic difference between the dynamic graph model used in [16] and TVG is as follows. In [16], the footprint is a clique and for TVG’s the footprint can be graphs with less number of edges as well. Moreover, after edge deletion/addition, port numbers are adjusted in the model of [16], which is not the case in the TVG model. This allows us to understand missing edges in the presence of 1-hop visibility. That’s why TVG exhibits a stronger graph model than the dynamic graph model used in [16]. The authors [16] ask the following question on the behaviour of the dispersion problem on TVG. Since the graph model of TVG is stronger than the dynamic model used in [16], their impossibility results do not hold here on TVGs.
In this work, we study our $k$-Dispersion on TVG. In the preliminary version of this work, we demonstrated that 1-hop visibility and global communication are necessary for solving $n$-Dispersion on TVG, and we provided a high-level approach for addressing $(n+1)$-Dispersion on TVG. Here, we extend those results and provide our algorithms in detail.
1.3 Our contribution
In this paper, we provide the following results.
  • Considering 1-hop visibility but without global communication, it is impossible to solve $k$-Dispersion on TVG when $k=pn$, $p\in\mathbb{N}$. This impossibility is valid even if the agents start from a rooted configuration, have knowledge of $n$ and are equipped with infinite memory (refer Section 2).
  • Considering global communication but without 1-hop visibility, it is impossible to solve $k$-Dispersion on TVG when $k=pn$, $p\in\mathbb{N}$. This impossibility is valid even if the agents start from a rooted configuration, have knowledge of $n$ and are equipped with infinite memory (refer Section 2).
  • We provide $\Omega(\log n)$ memory lower bound per agent to solve $k$-Dispersion on TVG (refer to Section 2), when $k>n$.
  • We present an algorithm that solves $(n+1)$-Dispersion on TVG when $n+1$ agents are initially co-located. This algorithm takes $O(n^{4}\cdot\log n)$ rounds and requires $O(\log n)$ memory per agent, considering local communication and zero visibility. This algorithm is memory-optimal (refer Section 3). If agents know $n$ and, $pn+1$ agents are co-located, then the same algorithm can be extended to solve $(pn+1)$-Dispersion considering local communication and zero visibility, where $p\in\mathbb{N}$ (refer to Section 4.1).
  • To solve $k$-Dispersion on TVG, we can use the existing algorithm of [16] when agents are equipped with 1-hop visibility and global communication. Further, the same algorithm of [16] with minor modifications works even when the adversary can delete any number of edges while keeping $\mathcal{G}$ connected in each round and $k(\geq 1)$ agents start from arbitrary initial configuration (refer to Section 4.2).
Figure 1: (a) If the adversary does not delete any edge at the beginning of round $r$, then this shows the computation of the agents. (b) If the adversary deletes $(u,\,u_{2})$ at the beginning of round $r$, then the view of agents at node $u_{1}$ does not change, but the view of agents at node $u$ and $u_{1}$ changes due to 1-hop visibility. In this case, agents at node $u$ can change $x_{1}$ to $x_{11}$, and agents at node $u_{2}$ can change $y_{3}$ to $y_{33}$ but agents at node $u_{1}$ do not change their decision because they can not understand deletion of edge $(u,\,u_{2})$ due to unavailability of global communication. Based on the cases mentioned in Theorem 2.1, the adversary decides which edge to delete at the start of round $r$. This outlines the high-level idea behind our proof Theorem 2.1.
2 Impossibility Results and Lower Bound
In this section, we show that it is impossible to solve $k$-Dispersion on TVG deterministically in our model if either global communication or 1-hop visibility is omitted, even if agents start from a rooted configuration. These impossibility results also hold if agents have infinite memory and knowledge of $n$. Recall that according to our model, the agents at a node $v$ cannot know whether an incident edge to node $v$ is missing or not unless agents are equipped with $1$-hop visibility.
Theorem 2.1

### v3

# Ashish Saxena Tanvir Kaur Kaushik MondaliiiiiiCorresponding Author

@@ Abstract
We aim to connect two problems, namely, dispersion and load balancing. Both problems have already been studied over static as well as dynamic graphs. Though dispersion and load balancing share some common features, the tools used in solving load balancing differ significantly from those used in solving dispersion. One of the reasons is that the load balancing problem is introduced and studied heavily over graphs where nodes are the processors and work under the message passing model, whereas dispersion is a task for mobile agents to achieve on graphs. To bring the (load) balancing aspect in the dispersion problem, we say, mobile agents move to balance themselves as equally as possible across the nodes of the graph, instead of stationary nodes sharing loads in the load balancing problem. We call it the $k$-balanced dispersion problem and study it on dynamic graphs. This is equivalent to the load balancing problem considering movable loads in form of the agents.
Earlier, on static graphs, the $k$-dispersion problem [TAMC 2019] aimed for the same by putting an upper bound on the number of agents on each node in the final configuration; however, the absence of a lower bound on the number of agents in their problem definition hampers the load-balancing aspect, as some nodes may end up with no agents in the final configuration. We take care of this part in our $k$-balanced dispersion problem definition and thus produce a stronger connection between the two domains.
We study the $k$-balanced dispersion problem on temporally connected dynamic graphs [TCS 2016] and $\ell$-bounded 1-interval connected dynamic graphs [JCSS 2021]. We provide several necessary and sufficient conditions to solve $k$-balanced dispersion problem including a couple of impossibility results.
@@
Mobile agents, Dispersion, Load balancing , Dynamic graphs, Deterministic algorithm
journal: Theoretical Computer Sciencejournal: NA
[…]
[1]organization=Department of Mathematics,addressline=Indian Institute of Technology Ropar, city=Rupnagar, postcode=140001, state=Punjab, country=India
@@ 1 Introduction
The dispersion problem was first introduced by Augustine et al. in [2] for static graphs. The goal is to coordinate $k\leq n$ mobile agents on a graph with $n$ nodes so that no two agents occupy the same node. In dispersion, the objective is to minimize the total cost of reaching such a configuration, assuming that movement is relatively cheap compared to the cost of multiple agents sharing the same node. This problem has been explored under a wide range of assumptions: different agent interaction mechanisms like whiteboards, tokens, face-to-face meetings, or global communication; different timing models such as asynchronous, semi-synchronous, or fully synchronous rounds; and varying degrees of knowledge about the graph and memory limitations. For a broader perspective, see [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Dispersion is closely related to several classical problems. Scattering [16, 17, 18] on graphs requires that $k\leq n$ agents spread out with equal spacing on symmetric structures such as rings or grids. This can be seen as dispersion with an additional constraint of equidistance. In multi-agent exploration [19], the goal is for agents starting from a common node to collectively visit all nodes in the graph as quickly as possible, which is a task that any dispersion algorithm with $k=n$ agents inherently solves. On the other hand, load balancing [20, 21] deals with nodes exchanging load to reach an even distribution. Given a graph $G$ representing the network with $n$ nodes, where each node contains the load $w_{i}$, the goal is to move the load across the edges so that finally the weight of each node is (approximately) equal to
$\overline{w_{i}}=\frac{1}{n}\sum_{j=1}^{n}w_{j}.$
Figure 1: Configuration as per $k$-dispersion.
Dispersion flips this perspective: instead of stationary nodes sharing load, mobile agents move to balance themselves across the network. While the tools used in load balancing differ significantly from those in mobile agent problems, studying dispersion can offer a common ground to connect these approaches and possibly transfer insights between the two domains. In 2019, Molla et al.[8] established a connection between the problems of dispersion and load balancing by introducing the notion of $k$-dispersion. Given a graph $G$ with $n$ nodes and $k\geq 1$ agents placed arbitrarily on its nodes, a configuration is said to be a $k$-dispersion if no node have more than $\lceil k/n\rceil$ agents. However, $k$-dispersion does not capture the idea of load balancing well enough. For example, consider $n=4$, $k=9$. Fig. 1 shows a valid configuration for $k$-dispersion as $\lceil k/n\rceil=3$. However, there is a node $v_{1}$ which has no agent. If parameters $k$ and $n$ are significantly large, there can be multiple nodes which has no agent. Therefore, agent load is not balanced. To build a stronger bridge between dispersion and load balancing, we introduce a constraint requiring each node to have at least $\lfloor k/n\rfloor$ and at most $\lceil k/n\rceil$ agents. Formally,
Definition 1.
($k$-balanced dispersion) Let $k\geq 1$ agents be placed on a graph $G$ with $n$ nodes. A configuration is said to be a balanced dispersion if each node has at least $\lfloor k/n\rfloor$ and at most $\lceil k/n\rceil$ agents.
In modern networks, such as mobile ad-hoc networks, network changes are frequent. These systems are often referred to as dynamic graphs. To analyze such systems, many dynamic graph models have been proposed, each under different assumptions. In the synchronous setting, where time is considered in discrete rounds, a dynamic graph $\mathcal{G}$ can be seen as a sequence of static graphs $\mathcal{G}_{0},\mathcal{G}_{1},\mathcal{G}_{2},\ldots$, where each $\mathcal{G}_{r}$ represents a graph at round $r$. This sequence is often referred to as a time evolving graph or dynamic graph. But if the graph evolves arbitrarily without any constraint, most distributed problems become unsolvable. To overcome this, several works introduce structural assumptions such as temporal connectivity [22], $T$-interval connectivity [23], $T$-path connectivity [24], or connectivity time [25], which have enabled progress on fundamental problems like exploration, gathering, gossiping, and leader election.
A graph $G$ is called port-labeled when each edge incident to a node $u$ is assigned a locally unique port number from the set $\{0,1,\ldots,\deg(u)-1\}$, where $\deg(u)$ is the degree of node $u$. That is, every undirected edge $\overline{uv}$ has two independent port labels: one at node $u$ and one at node $v$. This port numbering is crucial for enabling agents to distinguish between their outgoing edges. Without port labels, especially in anonymous graphs where nodes have no IDs, all edges at a node would be indistinguishable. As shown in [26], this makes even basic problems like exploration, dispersion, or rendezvous impossible to solve deterministically.
The question of how port numbers behave across rounds becomes even more critical in dynamic graphs. In the literature on dynamic graphs involving mobile agents, mainly two port-labeling models have been studied. The first, we call as fixed port labeling [27, 28, 29, 30, 31, 32], the network has a static footprint graph $G$, and each $\mathcal{G}_{r}$ is a subgraph of $G$. The outgoing ports at node $u$ in $\mathcal{G}_{r}$ are the same as the outgoing ports of node $u$ in $G$. Specifically, if port $\lambda$ at node $u$ leads to node $v$ in the footprint $G$, then whenever that edge corresponding port appears in any $\mathcal{G}_{r}$, it always leads to node $v$. This provides local stability in the network. The second, we call degree-based dynamic port labelling [33, 24], where there is no footprint. Therefore, the degree of each node can change arbitrarily in each round, and port numbers are reassigned dynamically based on the current degree. As a result, the port $\lambda$ at node $u$ may lead to different nodes at different rounds.
In the next section, we formally describe the model, and on this model, we study $k$-balanced dispersion.
2 Model
Dynamic graph: A dynamic network is modeled as a time-varying graph (TVG), denoted by $\mathcal{G}=(V,E,T,\rho)$, where $V$ is a set of nodes, $E$ is a set of edges, $T$ is the temporal domain, and $\rho:E\times T\rightarrow\{0,1\}$ is the presence function, which indicates whether a given edge is available at a given time. The static graph $G=(V,E)$ is referred to as the underlying graph (or footprint) of the TVG $\mathcal{G}$, where $|V|=n$ and $|E|=m$. For a node $v\in V$, let $E(v)\subseteq E$ denote the set of edges incident on $v$ in the footprint. The degree of node $v$ is defined as $\deg(v)=|E(v)|$, and the diameter of $\mathcal{G}_{r}$ is denoted by $D_{r}$, and $\widetilde{D}=\max_{r\geq 1}D_{r}$ denotes the dynamic diameter of $\mathcal{G}$. Nodes in $V$ are anonymous: they lack unique identifiers and possess no storage. Each edge incident to a node $v$ is locally labeled with a port number. This labeling is defined by a bijective function $\lambda_{v}:E(v)\rightarrow\{0,\ldots,\delta_{v}-1\}$, which assigns a distinct label to each incident edge of $v$. No further assumptions are made about the labeling. Assuming time is discrete, the TVG $\mathcal{G}$ can be viewed as a sequence of static graphs $\mathcal{S}_{\mathcal{G}}=\mathcal{G}_{0},\mathcal{G}_{1},\ldots,\mathcal{G}_{r},\ldots$, where each $\mathcal{G}_{r}=(V,E_{r})$ denotes the snapshot of $\mathcal{G}$ at round $r$, with $E_{r}=\{e\in E\mid\rho(e,r)=1\}$. The set of edges not present at time $r$ is denoted by $\overline{E}_{r}=E\setminus E_{r}\subseteq E$.
Dynamic graphs can be classified based on how their topological changes affect connectivity. The weakest meaningful requirement is temporal connectivity, which was introduced by Michail et al. [22] in 2016. Consider a set $\mathcal{J}=\{(\overline{u_{1}u_{2}},r_{1}),(\overline{u_{2}u_{3}},r_{2}),\ldots,(\overline{u_{k}u_{k+1}},r_{k})\}$, where $\{\overline{u_{1}u_{2}},~\overline{u_{2}u_{3}},~...,~\overline{u_{k}u_{k+1}}\}$ is a walk in $G$ (footprint) from $u_{1}$ to $u_{k+1}$ starting at round $r_{1}$ where $\forall\,i$, $\overline{u_{i}u_{i+1}}\in E_{i}$ and $r_{i+1}>r_{i}$. The set $\mathcal{J}$ is called a journey from $u_{1}$ to $u_{k+1}$. All possible journeys from node $u$ to $v$ starting on or after round $r$ are denoted by $\mathcal{J}(u,v,r)$. Formally,
Definition 2.
(temporal connectivity) [22] A dynamic graph $\mathcal{G}$ is said to be temporally connected (or connected over time) if for all $r\in\mathbb{N}\cup\{0\}$ and for all pair of nodes $u,v\in V$, there exists a journey from node $u$ to node $v$ starting at or after time $r$. In other words, $\mathcal{J}(u,v,r)\neq\emptyset$.iiiiiiiiiWe refer to a dynamic graph that satisfies temporal connectivity as a temporally connected graph.
Temporal connectivity is the minimal requirement for solving any global task in dynamic graphs, regardless of the initial positions of agents. In particular, any problem that requires every node to be involved, such as exploration or dispersion, is trivially unsolvable if the dynamic graph is not temporally connected. Stronger assumptions have also been studied in the literature. One well-known class of dynamic graphs guarantees connectivity at every round, rather than over time. A commonly used restriction is 1-interval connectivity, and a further refinement is its bounded variant.
Definition 3.
($\ell$-bounded 1-interval connectivity) [29] A dynamic graph $\mathcal{G}$ is 1-interval connected (or always connected) if every snapshot $\mathcal{G}_{r}\in\mathcal{S}_{\mathcal{G}}$ is connected. Furthermore, $\mathcal{G}$ is said to be $\ell$-bounded 1-interval connected if it is always connected and $|\overline{E}_{r}|$ is at most $\ell$.
@@ 2 Model
Agent: We consider $k$ agents to be present initially at nodes of the graph $G$. Each agent has a unique identifier assigned from the range $1,\,n^{c}]$, where $c$ is a constant. Each agent knows its ID but is unaware of the other agents’ IDs. Agents are not aware of the values of $n$, $k$, or $c$ unless stated otherwise. The agents are equipped with memory. An agent residing at a node, say $v$, in round $r$ knows $\deg(v)$ in the footprint $G$ and all associated ports corresponding to node $v$ in $G$; however, the agent does not understand if any incident edge of node $v$ is missing in $\mathcal{G}$ in round $t$. To be more precise, agent $a_{i}$ currently at node $v$ at round $r$ does not know the value of $\rho(e_{v},r)$, where $e_{v}$ is an edge incident at node $v$. Such a model has been considered in [29].
We call a node $v\in\mathcal{G}_{r}$ a hole at round $r$ if it has no agent, and a multinode if it has two or more agents.
Communication model: In this work, we consider an $l$-hop communication model, where $l\in\mathbb{N}\cup\{0\}$. Under this model, an agent at a node $v$ can send messages to all agents located within $l$ hops from $v$. When $l=0$, this reduces to face-to-face (f-2-f) communication [2, 3], meaning agents can only communicate if they are co-located at the same node. At the other hand, when $\widetilde{D}, ll-$hop communication becomes global communication [33, 34], allowing agents to exchange messages regardless of their positions in the network.
Visibility model: In this work, we use two types of visibility models: 0-hop visibility and 1-hop visibility. In the 0-hop visibility model [2, 4, 35], an agent at a node $v\in\mathcal{G}$ can see the IDs of agents present at $v$ in round $r$, as well as the port numbers at $v$, but nothing beyond that. In contrast, in the 1-hop visibility model [36, 37], an agent $a_{i}$ at node $v$ can also see all neighbors of $v$, including the IDs of agents (if any) at those neighboring nodes. Let $e_{v}$ be an edge incident to node $v$. In the 0-hop visibility model, agents cannot determine the value of $\rho(e_{v},r)$ at the beginning of round $r$. In contrast, under the 1-hop visibility model, they can determine this value at the beginning of round $r$.
@@
The algorithm runs in synchronous rounds. In each round $r$, an agent $a_{i}$ performs one Communicate-Compute-Move (CCM) cycle as follows:
  • 1.
@@
    Communicate: Agent $a_{i}$ at node $v_{i}$ can communicate with other agents $a_{j}$ present at the same node $v_{i}$ or present at any different node $v_{j}$, depending on the communication (f-2-f or global). The agent also understands whether it had a successful or an unsuccessful move in the last round.
  • 2.
@@
In this work, we use two terms to describe agents’ configurations within each round $r$: (i) at the beginning of round $r$, which refers to the agents’ configuration before the CCM cycle starts, and (ii) at the end of round $r$, which refers to the agents’ configuration after the CCM cycle completes.
If $k\geq 1$, and every node has at least $\lfloor k/n\rfloor$ and at most $\lceil k/n\rceil$ agents, then we refer to such a configuration as a dispersed configuration in this work. Otherwise, we refer to it as an undispersed configuration.
3 Related work
The problem of dispersion for $k\leq n$ agents has been extensively studied in static graphs under two main communication models: (a) f-2-f communication, and (b) global communication. In f-2-f, there is a wide range of literature [2, 3, 4, 5, 6, 7, 8, 9, 11, 12, 14, 15]. In 2022, Kshemkalyani et al. [13] studied the dispersion problem for $k\leq n$ when agents are equipped with global communication. A natural and fundamental question is whether such communication capabilities are strictly necessary to achieve dispersion. In 2024, Gorain et al. [10] partially addressed this question. They presented an algorithm for $k$ co-located agents that achieves dispersion without relying on f-2-f communication. In their model, the information available to an agent at a node $v$ at the beginning of each round is extremely limited. Specifically, each agent knows only the answers to the following two questions: (1) am I alone at node $v$? and (2) compared to the previous round, has the number of agents at $v$ increased or decreased? However, the question is still open: what is the necessary communication type to solve dispersion for $k\leq n$ agents?
Recently, researchers have shown growing interest in distributed problems such as exploration, gathering, dispersion, and leader election in dynamic graphs. This is motivated by the fact that dynamic graphs are more related to real-world networks, where topologies change over time. While dispersion has been well studied in static settings, relatively fewer works study the problem for $k\leq n$ agents in dynamic graphs [24, 33, 36]. Agarwalla et al. [36] consider a 1-bounded 1-interval connected dynamic ring and provide several deterministic algorithms. In contrast, Kshemkalyani et al. [33] and Saxena et al. [24] study dispersion in a more general and weaker dynamic model where the graph remains connected in every round, but there is no fixed footprint, and the adversary can arbitrarily add or remove edges.
In [33], two communication assumptions are considered: global communication and 1-neighborhood knowledge. This 1-neighbourhood knowledge is equivalent to saying 1-hop visibility. They proved the following two theorems:
Theorem 3.1.